# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@ 5.3. Nonabelian noncommutative Hodge Structure
$H^{1}_{Dol}(X{\setminus}0)$ carries a nonabelian mixed Hodge Structure.
As in the previous subsection, we consider the smoothing one-parameter family
$\{X_{t}\to X_{0}\},$
with the monodromy operator (see Fig. 9). Associated with the family is the spectrum of singularity, denoted by $S_{\alpha}$, encoding the Hodge-theoretic information.
@@
Figure 9. The monodromy of an isolated surface singularity.
Many of the classical aspects of Hodge theory have not been developed for the mixed nonabelian Hodge structures. In particular, no Clemens–Schmid sequence is known for mixed nonabelian Hodge structures. We propose the nonabelian noncommutative Hodge structure as a possible approach to Clemens–Schmid sequence for mixed nonabelian Hodge structures based on the theory of Landau–Ginzburg models. Our considerations suggest the following nonabelian noncommutative Hodge Structure, illustrated in Fig. 10.
[…]
Figure 10. Nonabelian noncommutative structure combining the moduli spaces of representations with Landau-Ginzburg theory.
The critical sets of the function $\varphi$ are the moduli spaces of connections, carrying the nonabelian Hodge structures. The proposals for constructing $\varphi$ using a reduction of the Chern–Simons were given, e.g. [21, 20]. The noncommutative Hodge structure is encoded in the global structure of $\varphi$ and in the spectrum, which labels the components, and the multiplicity corresponds to the Euler characteristics of the components, as explained in the previous subsection. The nonabelian noncommutative Hodge structure $\mathcal{H}$ can be thought of as a combination of these two phenomena. The properties of $\mathcal{H}$ are expected to be combinations of A and B sides of properties [39]. We formulate these as questions.
Question 5.4.
@@
    Does $\mathcal{H}$ have all properties of spectra, e.g. Thom–Sebastiani?
@@
    Does $\mathcal{H}_{spec}$ have Thom–Sebastiani and semi-continuity properties?
@@
    Does the Hitchin base $B$ of the moduli space $M_{Dol}(X{\setminus}0)$ embed into the space of stability conditions for Fukaya–Seidel category associated to the LG model $\mathcal{H}_{X}$?
    $\mathop{Stab}\mathop{FS}(\mathcal{H}(X))\supset B=\bigoplus\limits_{\lambda_{i}\in Spec}H_{\lambda_{i}}(Sym^{2}(X{\setminus}0))?$
@@
$X:x^{2}+y^{3}+z^{7}=0\supset$ Brieskorn homology sphere $\Sigma(2,3,7)$
@@
There are three irreducible flat $SL(2,\mathbb{C})$-connections, corresponding to three $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ orbits of the spectrum of $X$ given by
(5.11) $\frac{d_{1}}{2}+\frac{d_{2}}{3}+\frac{d_{3}}{7}$
with $d_{1}=1$, $d_{2}=1,2$ and $d_{3}=1,2,3,4,5,6$. The Landau–Ginzburg model for the link of $X$ was proposed in [20]. As pointed out in loc.cit., the Landau–Ginzburg potential may have additional critical points that have interpretation (and play an important role) in curve counting and also in 3d-3d correspondence [25, 26].
@@
Approach 1: In [1] Abouzaid-Manolescu show that the moduli space of $\mathop{SL}(2,\mathbb{C})$ representations of $\pi_{1}(Y)$ (for any 3-manifold $Y$) is a derived critical locus, and they prove it by choosing a Heegaard splitting. The question is how the space and the function $f$ such that $\text{crit}(f)=X$ depends on the choice of Heegaard splitting.
@@
Approach 2: Use the approach of [21, 20] to construct the potential. We must find a Lagrangian submanifold in a 3d Calabi–Yau manifold $T^{*}Y$ associated with $Y$; the disk counting then produces a potential $f$. Carrying this out might give a simpler proof of the above conjecture. Furthermore, the Fukaya-Seidel category for $f$ in this approach is expected to be related to the “category of line operators” $\text{MTC}[Y,G_{\mathbb{C}}]$ in [35, 8].
The spectra defined in the first and second approach may be the same. It seems conceivable that the following conjecture holds:
@@
    Hypercohomologies of the perverse sheaf of vanishing cycles $F$. When $F$ comes from Heegaard splitting of a 3-manifold $Y$, the above hypercohomologies form an invariant of $Y$ [1].
@@
    We can enhance the above hypercohomologies with a mixed Hodge structure. Combined with the monodromy, it leads to the derived spectrum. In the case of the derived spectrum connected with a moduli space of $\mathop{SL}(2,\mathbb{C})$ representations, it defines invariants of this moduli space and an invariant of the 3-manifold $Y$.
@@
    We can bring additional data to the above potential $f$ – a divisor $D$ along which $f$ has a log behavior. This allows an additional spectral grading from the number of times a path coming from one component of the critical set goes around $D$ and ending on another component of the critical set.
@@
Some of these ideas have appeared before in the theory of singularity and category theory – see, e.g., [40, 23]. This will require a new theory of spectra in the case of shifted symplectic structures.
5.4. Relation to vertex algebras and invariants of complex surfaces
The nonconventional modular properties of $\widehat{Z}_{\sigma}(q)$-invariants perfectly fit those exhibited by characters of a logarithmic vertex algebra. In fact, the relation of the form
(5.12) $\widehat{Z}_{\sigma}(q)=\chi_{\sigma}(q)$
is one of the predictions of the so-called “3d-3d correspondence” in physics, and can be viewed as its mathematical incarnation. Here, $\chi_{\sigma}(q)$ is a character of a log-VOA labeled by 3-manifold $Y$ (and a choice of the root system, which is implicit throughout the paper).
For a plumbed 3-manifold $Y$, a specific choice of a plumbing graph corresponds to a choice of a 4-manifold $X$, with $Y=\partial X$ as its boundary ($X$ is given by 4d Dehn surgery). If the plumbing is negative definite, this 4-manifold can be equipped with a complex structure (as a resolution of a normal surface singularity) and we can study analytical invariants of the complex surfaces and compare them to topological invariants of underlying 4-manifold and to invariants of $Y$. Since various $q$-series invariants of $X$ were available for quite some time, it is natural to ask how they compare with $\widehat{Z}_{\sigma}(q)$ and $Z_{0}(q)$. The development of $\widehat{Z}_{\sigma}(Y;q)$ was, in fact, largely motivated [35] by the connection to the Vafa–Witten and Donaldson–Thomas invariants of $X$ which share many similarities with $\widehat{Z}_{\sigma}(Y;q)$. For example, both require a choice of an additional structure whose cutting-and-gluing is described by a ‘decorated’ version of TQFT, and both are related to characters of VOA’s labeled by the corresponding 3-manifolds or 4-manifolds.
To keep this discussion more concrete and to explore these relations deeper, let us consider a rather special class of spherical 3-manifolds of $ADE$ type, i.e. $Y=S^{3}/\Gamma$, where $\Gamma$ is a discrete subgroup of $SU(2)$. In type $A$, $\Gamma=\mathbb{Z}_{p}$ is simply a cyclic group and $Y=L(p,1)$ is a Lens space. Similarly, for $E_{8}$ we get a Poincaré sphere, and we refer to other spherical 3-manifolds in this family by their type, e.g. as “$E_{6}$ manifold” and so on.
@@ 5.4. Relation to vertex algebras and invariants of complex surfaces
For a discrete subgroup $\Gamma\subset SU(2)$ of ADE type, let $X_{\Gamma}$ be the corresponding ALE space given by a resolution of the $\mathbb{C}^{2}/\Gamma$ singularity. Then, for any
(5.13) $\rho\in\mathrm{Hom}\left(\Gamma,GL(N)\right)$
the generating series of Vafa–Witten (and Donaldson–Thomas) invariants counting rank-$N$ sheaves on $X_{\Gamma}$ is a character of the affine Lie algebra $\widehat{\mathfrak{g}}$ at level $N$:
(5.14) $Z_{VW}(X_{\Gamma};q,\rho)\;=\;Z_{DT}(X_{\Gamma};q,\rho)\;=\;\chi_{\rho}^{\widehat{\mathfrak{g}}_{N}}(q)$
@@
Note that all representations $\rho:\Gamma\to GL(N)$ can be conjugated to $U(N)$ representations. In particular, this means that for spherical manifolds $S^{3}/\Gamma$ the problem of enumerating $U(N)$ flat connections is equivalent to the problem of classifying $GL(N)$ flat connections. [^18] This simple fact plays an important role in relating infinite-dimensional algebras and modular data associated to 4-manifolds and to 3-manifolds [22]; in particular, for spherical 3-manifolds the set $\{\rho\}$ labels simple objects in $\text{MTC}[Y]$ [8]. For example, when $\Gamma$ is of type $A$, i.e. a cyclic group $\mathbb{Z}_{p}$, we can think of $\rho$ as a Young tableau that can fit in a rectangle of (of the size determined by $N$ and the order of $\Gamma$). The number of such $\rho$ is equal to $\frac{(N+p-1)!}{(p-1)!N!}$. We can also use the relation $U(N)=\frac{U(1)\times SU(N)}{\mathbb{Z}_{N}}$ to enumerate $SU(N)$ or $SL(N)$ representations
(5.15) $\rho:\Gamma\to SL(N)$
Their number, equal to $\frac{(N+p-1)!}{(N-1)!p!}$, is the number of representations of $\widehat{\mathfrak{su}(N)}_{p}$.
@@
From the expressions in Example 4.5, we can write $\widehat{Z}_{0}$ and $\widehat{Z}_{1}$ in terms of false theta-functions:
(5.16) $\widehat{Z}_{0}(E_{6})$ $=\frac{2q^{-1}}{(q)_{\infty}}-\frac{q^{-\frac{25}{24}}}{(q)_{\infty}}\left(\tilde{\Psi}^{(1)}_{6}+\tilde{\Psi}^{(5)}_{6}\right)=\frac{q^{-1}}{(q)_{\infty}}\left(1-q+q^{2}+q^{5}-q^{7}+\ldots\right)$
$\widehat{Z}_{1}(E_{6})$ $=\frac{q^{-\frac{2}{3}-\frac{3}{8}}}{(q)_{\infty}}\Psi^{(3)}_{6}=\frac{q^{-2/3}}{(q)_{\infty}}\left(-1+q^{3}-q^{9}+q^{18}-q^{30}+q^{45}+\ldots\right)$
@@
where we restored the universal factor $\frac{1}{(q)_{\infty}}$ present for all 3-manifolds [^19] Due to its universal nature, this factor is often omitted for simplicity, but it is important in matching partition functions of 3d physical theory [34] and in comparing with characters of logarithmic vertex algebras [8, 9, 69] If we were to study higher rank invariants for more general root systems, there would be several such factors [57]. and used the standard notation for the false theta-functions
@@
(5.17) $\tilde{\Psi}^{(a)}_{p}(q)\;:=\;\sum_{n=0}^{\infty}\psi^{(a)}_{2p}(n)q^{\frac{n^{2}}{4p}}\qquad\in q^{\frac{a^{2}}{4p}}\,\mathbb{Z}[[q]],$
(5.20) $\psi^{(a)}_{2p}(n)=\left\{\begin{array}[]{cl}\pm 1,&n\equiv\pm a~{}\mod~{}2p\,,\\ 0,&\text{otherwise}.\end{array}\right.$
In particular, we have
(5.21) $\Delta_{\sigma}=h_{\sigma}-\frac{c}{24}=\{-1,\;-\tfrac{2}{3}\}$
which already gives some information about the central charge $c$ of the corresponding log-VOA and conformal weights $h_{\sigma}$ of its modules. In order to identify the right log-VOA and to match (5.16) with its characters, recall that $(p,p^{\prime})$ singlet log-VOA of type $A_{1}$ can be understood as a limit $p,q\to\infty$ of the $(p,q)$ Virasoro minimal model with the ratio $\frac{p}{q}=\frac{p}{p^{\prime}}$ held fixed. In turn, the $(p,q)$ Virasoro minimal model has $c=1-\frac{6(p-q)^{2}}{pq}$ and
(5.22) $h_{r,s}=\frac{(pr-qs)^{2}-(p-q)^{2}}{4pq}\,,\quad 1\leq r\leq q-1\,,\quad 1\leq s\leq p-1$
so that the minimal value of $h_{p,q}$ is $h_{\text{min}}=\frac{1-(p-q)^{2}}{4pq}$. A small calculation quickly shows that in the limit the central charge remains the same, whereas the value of $h_{\text{min}}$ is slightly modified:
(5.23) $c=1-\frac{6(p-p^{\prime})^{2}}{pp^{\prime}}\quad,\quad h_{\text{min}}=-\frac{(p-p^{\prime})^{2}}{4pp^{\prime}}$
so that
$c_{\text{eff}}=c-24h_{\text{min}}=1$
for all values of $p$ and $p^{\prime}$. The character of atypical modules of a $(p,p^{\prime})$ singlet log-VOA are
$\chi_{M_{r,s}}=\frac{1}{\eta(q)}\sum_{n\geq 0}\left(q^{\frac{1}{4pp^{\prime}}(2pp^{\prime}n+p^{\prime}r-ps)^{2}}-q^{\frac{1}{4pp^{\prime}}(2pp^{\prime}n+p^{\prime}r+ps)^{2}}\right)$
where $1\leq r\leq p^{\prime}$ and $1\leq s\leq p$. In particular, for a $(1,p)$ singlet log-VOA, we can write them in terms of false theta-functions (5.17) as
$\chi_{M_{1,s}}=\frac{1}{\eta(q)}\tilde{\Psi}_{p}^{(p-s)}$
@@
For example, from (5.16) we read off the value of $p=6$, which gives $c=13-6p-6p^{-1}=-24$. This, in turn, implies that $h_{\sigma}=\Delta_{\sigma}+\frac{c}{24}=\frac{(p-s)^{2}}{4p}+\frac{c-1}{24}=\{-1,-\frac{2}{3}\}$ for the $E_{6}$ manifold. Therefore, we conclude that for the $E_{6}$ manifold, $\widehat{Z}_{\sigma}(Y;q)$ are characters of a logarithmic $(1,6)$ singlet model.
Similarly, one can prove the following analog of Theorem 5.8:
@@
For the family of spherical 3-manifolds, $Y=S^{3}/\Gamma$, the non-perturbative $\mathop{SL}(2,\mathbb{C})$ invariants $\widehat{Z}_{\sigma}(Y;q)$ are equal to characters of the following (logarithmic) vertex algebras:
  • In type A, $\widehat{Z}_{\sigma}(Y;q)$ is a character of an ordinary (non-logarithmic) vertex algebra, namely the character of a Feigin-Fuchs module of the Virasoro algebra (or, equivalently, that of a $c=1$ “free boson” VOA).
  • In type Dn, $\widehat{Z}_{\sigma}(Y;q)$ is a character of a logarithmic $(1,p)$ singlet model, with $p=n-3$.
@@
In type $A$, the assertion follows from the explicit expression (see e.g. [35]):
(5.24) $\widehat{Z}_{\sigma}(Y;q)=\frac{1}{(q)_{\infty}}q^{\Delta_{\sigma}}$

### v3

@@ 5.2 Nonabelian noncommutative Hodge structure (NCNA)
$H^{1}_{\rm Dol}(X{\setminus}0)$ carries a nonabelian mixed Hodge structure.
As in the previous subsection, we consider the smoothing one-parameter family $\{X_{t}\to X_{0}\}$, with the monodromy operator (see Figure 9). Associated with the family is the spectrum of singularity, denoted by $S_{\alpha}$, encoding the Hodge-theoretic information.
$X_{t}$$X_{0}$
Figure 9: The monodromy of an isolated surface singularity.
Many of the classical aspects of Hodge theory have not been developed for mixed nonabelian Hodge structures. In particular, no Clemens–Schmid sequence is known for mixed nonabelian Hodge structures. We propose the nonabelian noncommutative Hodge structure (NCNA) as a possible approach to the Clemens–Schmid sequence for mixed nonabelian Hodge structures based on the theory of Landau–Ginzburg models. Our considerations suggest the following nonabelian noncommutative Hodge Structure, illustrated in Figure 10.
@@
Figure 10: Nonabelian noncommutative structure combining the moduli spaces of representations with Landau–Ginzburg theory.
The critical sets of the function $\varphi$ are the moduli spaces of connections, carrying the nonabelian Hodge structures. The proposals for constructing $\varphi$ using a reduction of the Chern–Simons functional were given, e.g., [20, 21]. The noncommutative Hodge structure (NC) is encoded in the global structure of $\varphi$ and in the spectrum, which labels the components, and the multiplicity corresponds to the Euler characteristic of the components, as explained in the previous subsection. The noncommutative nonabelian Hodge structure $\mathcal{H}$ can be thought of as a combination of these two phenomena. The properties of $\mathcal{H}$ are expected to be combinations of A and B sides of properties [38]. We formulate these as questions.
Question 5.7.
@@
    Does $\mathcal{H}$ have all properties of spectra, e.g., Thom–Sebastiani?
@@
    Does $\mathcal{H}$ have Thom–Sebastiani and semi-continuity properties?
@@
    Does the Hitchin base $B$ of the moduli space $M_{\rm Dol}(X{\setminus}0)$ embed into the space of stability conditions for Fukaya–Seidel category associated to the LG model $\mathcal{H}_{X}$?
    $\operatorname{Stab}\mathop{FS}(\mathcal{H}(X))\supset B=\bigoplus\limits_{\lambda_{i}\in{\rm Spec}}H_{\lambda_{i}}\bigl{(}{\rm Sym}^{2}(X{\setminus}0)\bigr{)}?$
@@
$X\colon x^{2}+y^{3}+z^{7}=0\supset$ Brieskorn homology sphere $\Sigma(2,3,7)$.
@@
There are three irreducible flat $\operatorname{SL}_{2}(\mathbb{C})$-connections, corresponding to the three $(\mathbb{Z}_{2})^{2}$ orbits of the spectrum of $X$ given by $\frac{d_{1}}{2}+\frac{d_{2}}{3}+\frac{d_{3}}{7}$ with $d_{1}=1$, $d_{2}=1,2$ and $d_{3}=1,2,3,4,5,6$. The Landau–Ginzburg model for the link of $X$ was proposed in [20]. As pointed out in loc.cit., the Landau–Ginzburg potential may have additional critical points that have interpretation (and play an important role) in curve counting and also in 3d-3d correspondence [25, 26].
@@
Approach $1$: In [1], Abouzaid–Manolescu show that the moduli space of $\operatorname{SL}_{2}(\mathbb{C})$ representations of $\pi_{1}(Y)$ (for any 3-manifold $Y$) is a derived critical locus, and they prove it by choosing a Heegaard splitting. The question is how the space and the function $f$ such that $\text{crit}(f)=X$ depend on the choice of Heegaard splitting.
@@
Approach $2$: Use the approach of [20, 21] to construct the potential. We must find a Lagrangian submanifold in a 3d Calabi–Yau manifold $T^{*}Y$ associated with $Y$; the disk counting then produces a potential $f$. Carrying this out might give a simpler proof of the above conjecture. Furthermore, the Fukaya–Seidel category for $f$ in this approach is expected to be related to the “category of line operators” $\text{MTC}[Y,G_{\mathbb{C}}]$ in [8, 34].
The spectra defined in the first and second approaches may be the same. It seems conceivable that the following conjecture holds.
@@
    Hypercohomologies of the perverse sheaf of vanishing cycles $F$. When $F$ comes from the Heegaard splitting of a 3-manifold $Y$, the above hypercohomologies form an invariant of $Y$ [1].
@@
    We can enhance the above hypercohomologies with a mixed Hodge structure. Combined with the monodromy, it leads to the derived spectrum. In the case of the derived spectrum connected with a moduli space of $\operatorname{SL}_{2}(\mathbb{C})$ representations, it defines invariants of this moduli space and an invariant of the 3-manifold $Y$.
@@
    We can bring additional data to the above potential $f$ — a divisor $D$ along which $f$ has a log behavior. This allows an additional spectral grading from the number of times a path coming from one component of the critical set goes around $D$ and ending on another component of the critical set.
@@
Some of these ideas have appeared before in the singularity theory and category theory — see, e.g., [23, 39]. This will require a new theory of spectra in the case of shifted symplectic structures.
5.3 Relation to vertex algebras and invariants of complex surfaces
The nonconventional modular properties of $\widehat{Z}_{\sigma}(q)$-invariants perfectly fit those exhibited by characters of a logarithmic vertex algebra. In fact, the relation of the form $\widehat{Z}_{\sigma}(q)=\chi_{\sigma}(q)$ is one of the predictions of the so-called “3d-3d correspondence” in physics, and can be viewed as its mathematical incarnation. Here, $\chi_{\sigma}(q)$ is a character of a log-VOA labeled by 3-manifold $Y$ (and a choice of the root system, which is implicit throughout the paper).
For a plumbed 3-manifold $Y$, a specific choice of a plumbing graph corresponds to a choice of a 4-manifold $X$, with $Y=\partial X$ as its boundary ($X$ is given by 4d Dehn surgery). If the plumbing is negative definite, this 4-manifold can be equipped with a complex structure (as a resolution of a normal surface singularity) and we can study analytical invariants of the complex surfaces and compare them to topological invariants of the underlying 4-manifold and to invariants of $Y$. Since various $q$-series invariants of $X$ were available for quite some time, it is natural to ask how they compare with $\widehat{Z}_{\sigma}(q)$ and $Z_{0}(q)$. The development of $\widehat{Z}_{\sigma}(Y;q)$ was, in fact, largely motivated [34] by the connection to the Vafa–Witten and Donaldson–Thomas invariants of $X$ which share many similarities with $\widehat{Z}_{\sigma}(Y;q)$. For example, both require a choice of an additional structure whose cutting-and-gluing is described by a ‘decorated’ version of TQFT, and both are related to characters of VOAs labeled by the corresponding 3-manifolds or 4-manifolds.
To keep this discussion more concrete and to explore these relations more deeply, let us consider a rather special class of spherical 3-manifolds of $ADE$ type, i.e., $Y=S^{3}/\Gamma$, where $\Gamma$ is a discrete subgroup of $\operatorname{SU}(2)$. In type $A$, $\Gamma=\mathbb{Z}_{p}$ is simply a cyclic group and $Y=L(p,1)$ is a Lens space. Similarly, for $E_{8}$ we get a Poincaré sphere, and we refer to other spherical 3-manifolds in this family by their type, e.g., as $E_{6}$ manifold” and so on.
@@ 5.3 Relation to vertex algebras and invariants of complex surfaces
For a discrete subgroup $\Gamma\subset\operatorname{SU}(2)$ of ADE type, let $X_{\Gamma}$ be the corresponding ALE space given by a resolution of the $\mathbb{C}^{2}/\Gamma$ singularity. Then, for any
$\rho\in\operatorname{Hom}(\Gamma,{\rm GL}(N))$
the generating series of Vafa–Witten ((and Donaldson–Thomas)) invariants counting rank-$N$ sheaves on $X_{\Gamma}$ is a character of the affine Lie algebra $\widehat{\mathfrak{g}}$ at level NN
$Z_{VW}(X_{\Gamma};q,\rho)=Z_{DT}(X_{\Gamma};q,\rho)=\chi_{\rho}^{\widehat{\mathfrak{g}}_{N}}(q),$
@@
Note that all representations $\rho\colon\Gamma\to{\rm GL}(N)$ can be conjugated to ${\rm U}(N)$ representations. In particular, this means that for spherical manifolds $S^{3}/\Gamma$ the problem of enumerating ${\rm U}(N)$ flat connections is equivalent to the problem of classifying ${\rm GL}(N)$ flat connections. [^17] This simple fact plays an important role in relating infinite-dimensional algebras and modular data associated to 4-manifolds and to 3-manifolds [22]; in particular, for spherical 3-manifolds the set $\{\rho\}$ labels simple objects in $\text{MTC}[Y]$ [8]. For example, when $\Gamma$ is of type $A$, i.e., a cyclic group $\mathbb{Z}_{p}$, we can think of $\rho$ as a Young tableau that can fit in a rectangle (of the size determined by $N$ and the order of $\Gamma$). The number of such $\rho$ is equal to $\frac{(N+p-1)!}{(p-1)!N!}$. We can also use the relation ${\rm U}(N)=\frac{{\rm U}(1)\times\operatorname{SU}(N)}{\mathbb{Z}_{N}}$ to enumerate $\operatorname{SU}(N)$ or $\operatorname{SL}(N)$ representations $\rho\colon\Gamma\to\operatorname{SL}(N)$. Their number, equal to $\frac{(N+p-1)!}{(N-1)!p!}$, is the number of representations of $\widehat{\mathfrak{su}(N)}_{p}$.
@@
From the expressions in Example 4.5, we can write $\widehat{Z}_{0}$ and $\widehat{Z}_{1}$ in terms of false theta-functions
$\widehat{Z}_{0}(E_{6})$ $=\frac{2q^{-1}}{(q)_{\infty}}-\frac{q^{-\frac{25}{24}}}{(q)_{\infty}}\bigl{(}\tilde{\Psi}^{(1)}_{6}+\tilde{\Psi}^{(5)}_{6}\bigr{)}=\frac{q^{-1}}{(q)_{\infty}}\bigl{(}1-q+q^{2}+q^{5}-q^{7}+\cdots\bigr{)},$
$\widehat{Z}_{1}(E_{6})$ $=\frac{q^{-\frac{2}{3}-\frac{3}{8}}}{(q)_{\infty}}\tilde{\Psi}^{(3)}_{6}=\frac{q^{-2/3}}{(q)_{\infty}}\bigl{(}-1+q^{3}-q^{9}+q^{18}-q^{30}+q^{45}+\cdots\bigr{)},$ (5.11)
@@
where we restored the universal factor $\frac{1}{(q)_{\infty}}$ present for all 3-manifolds [^18] Due to its universal nature, this factor is often omitted for simplicity, but it is important in matching partition functions of 3d physical theory [33] and in comparing with characters of logarithmic vertex algebras [8, 9, 68]. If we were to study higher rank invariants for more general root systems, there would be several such factors [57]. and used the standard notation for the false theta-functions
@@
$\tilde{\Psi}^{(a)}_{p}(q):=\sum_{n=0}^{\infty}\psi^{(a)}_{2p}(n)q^{\frac{n^{2}}{4p}}\in q^{\frac{a^{2}}{4p}}\mathbb{Z}[[q]],\qquad\psi^{(a)}_{2p}(n)=\begin{cases}\pm 1,&n\equiv\pm a~\mod~2p,\\ 0,&\text{otherwise}.\end{cases}$ (5.12)
In particular, we have $\Delta_{\sigma}=h_{\sigma}-\frac{c}{24}\in\bigl{\{}-1,-\tfrac{2}{3}\bigr{\}}$, which already gives some information about the central charge $c$ of the corresponding log-VOA and conformal weights $h_{\sigma}$ of its modules. Let $a,b\in\mathbb{Z}_{\geq 1}$}, $\gcd(a,b)=1$. For $r^{\prime},s^{\prime}\in\mathbb{Z}$, set
$\beta_{r^{\prime},s^{\prime}}:=\left(-\frac{r^{\prime}-1}{a}+\frac{s^{\prime}-1}{b}\right)\sqrt{ab}\varpi,$
where $\varpi=\frac{1}{2}\alpha$ is the fundamental weight of $\mathfrak{g}=\mathfrak{sl}_{2}$. In particular, for $n,n_{1},n_{2}\in\mathbb{Z}$, $n_{1}+n_{2}=n$ and $1\leq s\leq a$, $1\leq r\leq b$, we have
$\beta(-n,s,r):=\left(-n-\frac{s-1}{a}+\frac{r-1}{b}\right)\sqrt{ab}\varpi=\beta_{2-r+bn_{1},2-s-an_{2}}$
and let us denote by $\pi_{\beta}$ the Fock space with weight $\beta$. For the conformal vector
$\omega:=\frac{1}{2}\alpha_{(-1)}\varpi+Q_{0}\varpi_{(-2)}|0\rangle\in\pi_{0},\qquad Q_{0}:=\sqrt{ab}\left(\frac{1}{a}-\frac{1}{b}\right)=\sqrt{\frac{b}{a}}-\sqrt{\frac{a}{b}},$
the central charge $c$ and the conformal weight $h_{r^{\prime},s^{\prime}}$ of $|\beta_{r^{\prime},s^{\prime}}\rangle\in\pi_{\beta_{r^{\prime},s^{\prime}}}$ is given by
$c=1-12|Q_{0}\varpi|^{2}=1-\frac{6(a-b)^{2}}{ab}=1-6\left(\frac{a}{b}-2+\frac{b}{a}\right),$
$h_{r^{\prime},s^{\prime}}=\frac{1}{2}|\beta_{r^{\prime},s^{\prime}}-Q_{0}\varpi|^{2}+\frac{c-1}{24}=\frac{(ar^{\prime}-bs^{\prime})^{2}-(a-b)^{2}}{4ab}$
[…]
Therefore, the character $\chi_{M}:=\text{tr}_{M}q^{L_{0}-\frac{c}{24}}$ of $\pi_{\beta_{r^{\prime},s^{\prime}}}$ is
$\chi_{\pi_{\beta_{r^{\prime},s^{\prime}}}}=\frac{1}{(q)_{\infty}}q^{h_{r^{\prime},s^{\prime}}-\frac{c}{24}}=\frac{1}{\eta(q)}q^{\frac{1}{2}|\beta_{r^{\prime},s^{\prime}}-Q_{0}\varpi|^{2}}.$
In particular,
$\chi_{\pi_{\beta(-n;s,r)}}=\frac{1}{\eta(q)}q^{\frac{1}{2}|\sqrt{ab}(-n-\frac{s}{a}+\frac{r}{b})\varpi|^{2}}=\frac{1}{\eta(q)}q^{\frac{1}{4ab}(-nab-bs+ar)^{2}}.$
For the case $(a,b)=(1,p)$ with $p\geq 2$ (so that $s=1$ and $1\leq r\leq p$), we have
$\chi_{\pi_{\beta(-n;1,r)}}=\frac{1}{\eta(q)}q^{\frac{1}{4p}(-(n+1)p+r)^{2}}.$
By the Felder complex, the character of the irreducible module $M(0;1,r)\subset\pi_{\beta(0;1,r)}$ of the $(1,p)$ singlet log-VOA $M(0;1,1)\subset\pi_{\beta(0;1,1)}$ is given by
$\chi_{M(0;1,r)}=\sum_{m\geq 0}\chi_{\pi_{\beta(-2m;1,r)}}-\chi_{\pi_{\beta(-(2m+1);1,p-r)}}=\frac{1}{\eta(q)}\sum_{m\geq 0}q^{\frac{1}{4p}(-(2m+1)p\pm r)^{2}}=\tilde{\Psi}_{p}^{(p-r)}.$
@@
For example, from (5.11) we read off the value of $p=6$, which gives $c=13-6p-6p^{-1}=-24$. This, in turn, implies that $h_{\sigma}=\Delta_{\sigma}+\frac{c}{24}=\frac{(p-s)^{2}}{4p}+\frac{c-1}{24}=\bigl{\{}-1,-\frac{2}{3}\bigr{\}}$ for the $E_{6}$ manifold. Therefore, we conclude that for the $E_{6}$ manifold, $\widehat{Z}_{\sigma}(Y;q)$ are characters of a logarithmic $(1,6)$ singlet model.
Similarly, one can prove the following analogue of Theorem 5.11.
@@
For the family of spherical $3$-manifolds, $Y=S^{3}/\Gamma$, the non-perturbative $\operatorname{SL}_{2}(\mathbb{C})$ invariants $\widehat{Z}_{\sigma}(Y;q)$ are equal to characters of the following ((logarithmic)) vertex algebras:
  • In type A, $\widehat{Z}_{\sigma}(Y;q)$ is a character of an ordinary ((non-logarithmic)) vertex algebra, namely the character of a Feigin–Fuchs module of the Virasoro algebra ((or, equivalently, that of a c=1{c=1} “free boson” VOA)).
  • In type Dn, $\widehat{Z}_{\sigma}(Y;q)$ is a character of a logarithmic $(1,p)$ singlet model, with p=n3p=n-3.
@@
In type $A$, the assertion follows from the explicit expression (see, e.g., [34])
$\widehat{Z}_{\sigma}(Y;q)=\frac{1}{(q)_{\infty}}q^{\Delta_{\sigma}}.$